# Latent Factor Analysis in Short Panels

### v2

@@ E.5.1 Asymptotic expansions
We first establish the asymptotic expansion of $\hat{\theta}$ along the lines of pseudo maximum likelihood estimators (White (1982)). The sample criterion is $\hat{L}(\theta)$ given in Equation (E.1), where $\theta=\left(vec(F)^{\prime},diag(V_{\varepsilon})^{\prime}\right)^{\prime}$ is subject to the nonlinear vector constraint $g(\theta):=\{[F^{\prime}V_{\varepsilon}^{-1}F]_{i,j}\}_{i<j}=0$, i.e., matrix $F^{\prime}V_{\varepsilon}^{-1}F$ is diagonal. By standard methods for constrained M-estimators, we consider the FOC of the Lagrangian function: $\frac{\partial\hat{L}(\hat{\theta})}{\partial\theta}-\frac{\partial g(\hat{\theta})^{\prime}}{\partial\theta}\hat{\lambda}_{L}=0$ and $g(\hat{\theta})=0$, where $\hat{\lambda}_{L}$ is the $\frac{1}{2}k(k-1)$ dimensional vector of estimated Lagrange multipliers. Define vector $\tilde{\theta}:=\left(vec(F_{0})^{\prime},diag(\tilde{V}_{\varepsilon})^{\prime}\right)^{\prime}$, which also satisfies the constraint $g(\tilde{\theta})=0$ by the in-sample factor normalization. We apply the mean value theorem to the FOC around $\tilde{\theta}$ and get:
$\hat{J}(\bar{\theta})\sqrt{n}(\hat{\theta}-\tilde{\theta})+A(\hat{\theta})\sqrt{n}\hat{\lambda}_{L}$ $=$ $\sqrt{n}\frac{\partial\hat{L}(\tilde{\theta})}{\partial\theta},$ (E.16)
$A(\bar{\theta})^{\prime}\sqrt{n}(\hat{\theta}-\tilde{\theta})$ $=$ $0,$ (E.17)
where $\hat{J}(\theta):=-\frac{\partial^{2}{\hat{L}}({\theta})}{\partial\theta\partial\theta^{\prime}}$ is the $r\times r$ Hessian matrix, $A(\theta):=\frac{\partial g({\theta})^{\prime}}{\partial\theta}$ is the $r\times\frac{1}{2}k(k-1)$ dimensional gradient matrix of the constraint function, and $\bar{\theta}$ is a mean value vector between $\hat{\theta}$ and $\tilde{\theta}$ componentwise. Matrix $A(\theta)$ is full rank for $\theta$ in a neighbourhood of $\theta_{0}$. For any $\theta$ define the $r\times(r-\frac{1}{2}k(k-1))$ matrix $B(\theta)$ with orthonormal columns that span the orthogonal complement of the range of $A(\theta)$. Matrix function $B(\theta)$ is continuous in $\theta$ in a neighbourhood of $\theta_{0}$. [^38] Matrix $B(\theta)$ is uniquely defined up to rotation and sign changes in their columns. We can pick a unique representer such that matrix $B(\theta)$ is locally continuous, e.g., by taking $B(\theta)=\tilde{B}(\theta)[\tilde{B}(\theta)^{\prime}\tilde{B}(\theta)]^{-1/2}$, where matrix $\tilde{B}(\theta)$ consists of the first $r-\frac{1}{2}k(k-1)$ columns of $I_{r}-A(\theta)[A(\theta)^{\prime}A(\theta)]^{-1}A(\theta)^{\prime}$, if those columns are linearly independent. Then, by multiplying Equation (E.16) times $B(\hat{\theta})^{\prime}$ to get rid of the Lagrange multiplier vector, using the identity $I_{r}=A(\theta)(A(\theta)^{\prime}A(\theta))^{-1}A(\theta)^{\prime}+B(\theta)B(\theta)^{\prime}$ for $\theta=\bar{\theta}$ and Equation (E.17), we get $[B(\hat{\theta})^{\prime}\hat{J}(\bar{\theta})B(\bar{\theta})]B(\bar{\theta})^{\prime}\sqrt{n}(\hat{\theta}-\tilde{\theta})=B(\hat{\theta})^{\prime}\sqrt{n}\frac{\partial\hat{L}(\tilde{\theta})}{\partial\theta}$. By the uniform convergence of $\hat{J}(\theta)$ to $J(\theta):=-\frac{\partial^{2}{L}_{0}({\theta})}{\partial\theta\partial\theta^{\prime}}$, and the consistency of the FA estimator $\hat{\theta}$ (Section E.2), matrix $B(\hat{\theta})^{\prime}\hat{J}(\bar{\theta})B(\bar{\theta})$ converges to $B_{0}^{\prime}J_{0}B_{0}$, where $J_{0}:=J(\theta_{0})$ and $B_{0}:=B(\theta_{0})$. Matrix $B_{0}^{\prime}J_{0}B_{0}$ is invertible under the local identification Assumption A.5 (see Lemma 6 condition d)). Then, $B(\bar{\theta})^{\prime}\sqrt{n}(\hat{\theta}-\tilde{\theta})=[B(\hat{\theta})^{\prime}\hat{J}(\bar{\theta})B(\bar{\theta})]^{-1}B(\hat{\theta})^{\prime}\sqrt{n}\frac{\partial\hat{L}(\tilde{\theta})}{\partial\theta}$ w.p.a. $1$. By using again $I_{r}=A(\bar{\theta})(A(\bar{\theta})^{\prime}A(\bar{\theta}))^{-1}A(\bar{\theta})^{\prime}+B(\bar{\theta})B(\bar{\theta})^{\prime}$ and Equation (E.17), we get $\sqrt{n}(\hat{\theta}-\tilde{\theta})=B(\bar{\theta})[B(\hat{\theta})^{\prime}\hat{J}(\bar{\theta})B(\bar{\theta})]^{-1}B(\hat{\theta})^{\prime}\sqrt{n}\frac{\partial\hat{L}(\tilde{\theta})}{\partial\theta}$. The distributional results established below imply $\sqrt{n}\frac{\partial\hat{L}(\tilde{\theta})}{\partial\theta}=O_{p}(1)$. Thus, we get $\sqrt{n}$-consistency:
$\sqrt{n}(\hat{\theta}-\tilde{\theta})=B_{0}(B_{0}^{\prime}J_{0}B_{0})^{-1}B_{0}^{\prime}\sqrt{n}\frac{\partial\hat{L}(\tilde{\theta})}{\partial\theta}+o_{p}(1).$ (E.18)
Let us now find the score $\frac{\partial\hat{L}({\theta})}{\partial\theta}$. We have $\frac{\partial\hat{L}({\theta})}{\partial\theta}=\left(\frac{\partial vec(\Sigma(\theta))}{\partial\theta^{\prime}}\right)^{\prime}vec\left(\frac{\partial\hat{L}(\theta)}{\partial\Sigma}\right),$ where $vec\left(\frac{\partial\hat{L}(\theta)}{\partial\Sigma}\right)=\frac{1}{2}\left(\Sigma(\theta)^{-1}\otimes\Sigma(\theta)^{-1}\right)vec\left(\hat{V}_{y}-\Sigma(\theta)\right).$ Moreover, by using $vec(\Sigma(\theta))=\sum_{j=1}^{k}F_{j}\otimes F_{j}+[e_{1}\otimes e_{1}~{}:~{}\cdots~{}:~{}e_{T}\otimes e_{T}]diag(V_{\varepsilon})$, where $e_{t}$ is the $t$-th column of $I_{T}$, we get: $\frac{\partial vec(\Sigma(\theta))}{\partial\theta^{\prime}}=\left[(I_{T}\otimes F_{1})+(F_{1}\otimes I_{T})~{}:~{}\cdots~{}:~{}(I_{T}\otimes F_{k})+(F_{k}\otimes I_{T})~{}:~{}\right.$ $\left.e_{1}\otimes e_{1}~{}:~{}\cdots~{}:~{}e_{T}\otimes e_{T}\right].$ Thus, we get: $\sqrt{n}\frac{\partial\hat{L}(\tilde{\theta})}{\partial\theta}=\frac{1}{2}\left(\frac{\partial vec(\Sigma(\tilde{\theta}))}{\partial\theta^{\prime}}\right)^{\prime}\left(\tilde{V}_{y}^{-1}\otimes\tilde{V}_{y}^{-1}\right)\sqrt{n}vec\left(\hat{V}_{y}-\tilde{V}_{y}\right)$. From Equation (E.2) and Lemma 5 we have $\hat{V}_{y}=\tilde{V}_{y}+\frac{1}{\sqrt{n}}(Z_{n}+W_{n}F^{\prime}+FW_{n}^{\prime})+o_{p}(\frac{1}{\sqrt{n}})$, where $W_{n}:=\frac{1}{\sqrt{n}}\varepsilon\beta$. Thus, $\sqrt{n}\frac{\partial\hat{L}(\tilde{\theta})}{\partial\theta}=\frac{1}{2}\left(\frac{\partial vec(\Sigma({\theta}_{0}))}{\partial\theta^{\prime}}\right)^{\prime}\left({V}_{y}^{-1}\otimes{V}_{y}^{-1}\right)vec\left(W_{n}F^{\prime}+FW_{n}^{\prime}+Z_{n}\right)+o_{p}(1)$ and, from Equation (E.18), we get:
@@
$\sqrt{n}(\hat{\theta}-\tilde{\theta})=\frac{1}{2}B_{0}\left(B_{0}^{\prime}J_{0}B_{0}\right)^{-1}B_{0}^{\prime}\left(\frac{\partial vec(\Sigma({\theta}_{0}))}{\partial\theta^{\prime}}\right)^{\prime}\left({V}_{y}^{-1}\otimes{V}_{y}^{-1}\right)vec\left(W_{n}F^{\prime}+FW_{n}^{\prime}+Z_{n}\right)+o_{p}(1).$ (E.19)
E.5.2 Asymptotic normality
@@
In this subsection, we establish the asymptotic normality of estimators $\hat{F}$ and $\hat{V}_{\varepsilon}$. From Lemma 1, as $n\rightarrow\infty$ and $T$ is fixed, we have the Gaussian distributional limit $Z_{n}\Rightarrow Z$ with $vech(Z)\sim N(0,\Omega_{Z})$, where the asymptotic variance $\Omega_{Z}$ is related to the asymptotic variance $\Omega$ of $\mathscr{Z}$ such that $Cov(Z_{ts},Z_{rp})=\sqrt{V_{\varepsilon,tt}V_{\varepsilon,ss}V_{\varepsilon,rr}V_{\varepsilon,pp}}Cov(\mathscr{Z}_{ts},\mathscr{Z}_{rp})$. Moreover, $Z_{n}^{*}\Rightarrow Z^{*}=G^{\prime}V_{\varepsilon}^{-1}ZV_{\varepsilon}^{-1}G$ and $\bar{Z}_{n}:=Z_{n}-\mathcal{T}_{F,V_{\varepsilon}}(Z_{n})\Rightarrow\bar{Z}$, where $\bar{Z}=Z-\mathcal{T}_{F,V_{\varepsilon}}(Z)=Z-V_{\varepsilon}^{2}diag\left((\boldsymbol{X}^{\prime}\boldsymbol{X})^{-1}\boldsymbol{X}^{\prime}vech(Z^{*})\right)$ (see (E.7)). The distributional limit of $W_{n}$ is given next.
@@
Under Assumptions 1, 2 and A.2, A.3, A.8, as $n\rightarrow\infty$, (a) we have $W_{n}\Rightarrow\bar{W}$, where $vec(\bar{W})\sim N(0,\Omega_{W})$ with $\Omega_{W}=Q_{\beta}\otimes V_{\varepsilon}$, and (b) if additionally $E[w_{i,t}w_{i,r}w_{i,s}]=0$, for all $t,r,s$ and $i$, then $Z$ and $\bar{W}$ are independent.
We get the following proposition from Lemmas 1 and 7 (see proof at the end of the section).
@@
Under Assumptions 1-2 and A.1-A.6, A.8, as $n\rightarrow\infty$ and $T$ is fixed, for $j=1,..,k$:
$\sqrt{n}diag(\hat{V}_{\varepsilon}-\tilde{V}_{\varepsilon})\Rightarrow V_{\varepsilon}^{2}(\boldsymbol{X}^{\prime}\boldsymbol{X})^{-1}\boldsymbol{X}^{\prime}vech(Z^{*}),$ (E.20)
$\sqrt{n}(\hat{F}_{j}-F_{j})\Rightarrow R_{j}(\bar{W}F^{\prime}+F\bar{W}^{\prime}+\bar{Z})V_{\varepsilon}^{-1}F_{j}+\Lambda_{j}\{[V_{\varepsilon}(\boldsymbol{X}^{\prime}\boldsymbol{X})^{-1}\boldsymbol{X}^{\prime}vech({Z}^{*})]\odot F_{j}\},\qquad\quad$ (E.21)
$\sqrt{n}(\hat{F}_{j}\hat{\mathcal{D}}-F_{j})\Rightarrow\frac{1}{\gamma_{j}}(\bar{W}F^{\prime}+F\bar{W}^{\prime}+\bar{Z})V_{\varepsilon}^{-1}F_{j},\qquad\quad$ (E.22)
@@
The joint asymptotic Gaussian distribution of the FA estimators involves the Gaussian matrices $Z^{*}$, $\bar{Z}$ and $\bar{W}$, the former two being symmetric. The asymptotic distribution of $\hat{V}_{\varepsilon}$ involves recentering around $\tilde{V}_{\varepsilon}=\frac{1}{n}\sum_{i=1}^{n}E[\varepsilon_{i}\varepsilon_{i}^{\prime}]$, i.e., the finite-sample average cross-moments of errors, and not $V_{\varepsilon}$. For the asymptotic distribution of any functional that depends on $F$ up to one-to-one transformations of its columns, we can use the Gaussian law of (E.22) involving $\bar{W}$ and $\bar{Z}$ only.
The asymptotic expansions (E.20)-(E.21) characterize explicitly the matrices $C_{1}(\theta)$ and $C_{2}(\theta)$ that appear in Theorem 2 in Anderson and Amemiya (1988). Their derivation is based on an asymptotic normality argument treating $\hat{\theta}$ as a M-estimator, see Section C.2. However, neither the asymptotic variance nor a feasible CLT are given in Anderson and Amemiya (1988). We cannot use their results for our empirics.
@@
Finally, by applying the CLT to (E.19), the asymptotic distribution of vector $\hat{\theta}$ is:
$\sqrt{n}(\hat{\theta}-\tilde{\theta})\Rightarrow\frac{1}{2}B_{0}\left(B_{0}^{\prime}J_{0}B_{0}\right)^{-1}B_{0}^{\prime}\left(\frac{\partial vec(\Sigma({\theta}_{0}))}{\partial\theta^{\prime}}\right)^{\prime}\left({V}_{y}^{-1}\otimes{V}_{y}^{-1}\right)vec\left(\bar{W}F^{\prime}+F\bar{W}^{\prime}+Z\right).\quad$ (E.23)
The Gaussian asymptotic distribution in (E.23) matches those in (E.20) and (E.21) written for the components, and its asymptotic variance yields the ‘sandwich formula”. The result in (E.23) is analogue to Theorem 2 in Anderson and Amemiya (1988), for different factor normalization and recentering of the variance estimator.
Proof of Proposition 8: From (E.7), we have the asymptotic expansion: $\sqrt{n}diag(\hat{V}_{\varepsilon}-\tilde{V}_{\varepsilon})=diag(\Psi_{\varepsilon})+o_{p}(1)=V_{\varepsilon}^{2}(\boldsymbol{X}^{\prime}\boldsymbol{X})^{-1}\boldsymbol{X}^{\prime}vech(Z_{n}^{*})+o_{p}(1).$ Moreover, from Proposition 7 (a) and using $\Psi_{y}-\Psi_{\varepsilon}=W_{n}F^{\prime}+FW_{n}^{\prime}+\bar{Z}_{n}$, we have: $\sqrt{n}(\hat{F}_{j}-F_{j})=R_{j}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F_{j}+\Lambda_{j}\Psi_{\varepsilon}V_{\varepsilon}^{-1}F_{j}+o_{p}(1)=R_{j}(W_{n}F^{\prime}+FW_{n}^{\prime}+\bar{Z}_{n})V_{\varepsilon}^{-1}F_{j}+\Lambda_{j}[diag(\Psi_{\varepsilon})\odot(V_{\varepsilon}^{-1}F_{j})]+o_{p}(1)=R_{j}(W_{n}F^{\prime}+FW_{n}^{\prime}+\bar{Z}_{n})V_{\varepsilon}^{-1}F_{j}+\Lambda_{j}\{[V_{\varepsilon}(\boldsymbol{X}^{\prime}\boldsymbol{X})^{-1}\boldsymbol{X}^{\prime}vech(Z_{n}^{*})]\odot F_{j}\}+o_{p}(1).$ Lemmas 1 and 7 yield (E.20)-(E.21), together with (E.22) from (E.8) since $\Psi_{y}-\Psi_{\varepsilon}\Rightarrow\bar{W}F^{\prime}+F\bar{W}^{\prime}+\bar{Z}$.
E.5.3 Feasible CLT for the FA estimators

### v3

@@ B.5.1 Asymptotic expansions
We first establish the asymptotic expansion of $\hat{\theta}$ along the lines of pseudo maximum likelihood estimators (White (1982)). The sample criterion is $\hat{L}(\theta)$ given in Equation (B.1), where $\theta=\left(vec(F)^{\prime},diag(V_{\varepsilon})^{\prime}\right)^{\prime}$ is subject to the nonlinear vector constraint $g(\theta):=\{[F^{\prime}V_{\varepsilon}^{-1}F]_{i,j}\}_{i<j}=0$, i.e., matrix $F^{\prime}V_{\varepsilon}^{-1}F$ is diagonal. By standard methods for constrained M-estimators, we consider the FOC of the Lagrangian function: $\frac{\partial\hat{L}(\hat{\theta})}{\partial\theta}-\frac{\partial g(\hat{\theta})^{\prime}}{\partial\theta}\hat{\lambda}_{L}=0$ and $g(\hat{\theta})=0$, where $\hat{\lambda}_{L}$ is the $\frac{1}{2}k(k-1)$ dimensional vector of estimated Lagrange multipliers. Define vector $\tilde{\theta}:=\left(vec(F_{0})^{\prime},diag(\tilde{V}_{\varepsilon})^{\prime}\right)^{\prime}$, which also satisfies the constraint $g(\tilde{\theta})=0$ by the in-sample factor normalization. We apply the mean value theorem to the FOC around $\tilde{\theta}$ and get:
$\hat{J}(\bar{\theta})\sqrt{n}(\hat{\theta}-\tilde{\theta})+A(\hat{\theta})\sqrt{n}\hat{\lambda}_{L}$ $=$ $\sqrt{n}\frac{\partial\hat{L}(\tilde{\theta})}{\partial\theta},$ (B.16)
[…]
$\sqrt{n}(\hat{\theta}-\tilde{\theta})=B_{0}(B_{0}^{\prime}J_{0}B_{0})^{-1}B_{0}^{\prime}\sqrt{n}\frac{\partial\hat{L}(\tilde{\theta})}{\partial\theta}+o_{p}(1).$ (B.18)
[…]
@@
$\sqrt{n}(\hat{\theta}-\tilde{\theta})=\frac{1}{2}B_{0}\left(B_{0}^{\prime}J_{0}B_{0}\right)^{-1}B_{0}^{\prime}\left(\frac{\partial vec(\Sigma({\theta}_{0}))}{\partial\theta^{\prime}}\right)^{\prime}\left({V}_{y}^{-1}\otimes{V}_{y}^{-1}\right)vec\left(W_{n}F^{\prime}+FW_{n}^{\prime}+Z_{n}\right)+o_{p}(1).$ (B.19)
B.5.2 Asymptotic normality
@@
In this subsection, we establish the asymptotic normality of estimators $\hat{F}$ and $\hat{V}_{\varepsilon}$. From Lemma 2, as $n\rightarrow\infty$ and $T$ is fixed, we have the Gaussian distributional limit $Z_{n}\Rightarrow Z$ with $vech(Z)\sim N(0,\Omega_{Z})$, where the asymptotic variance $\Omega_{Z}$ is related to the asymptotic variance $\Omega$ of $\mathscr{Z}$ such that $Cov(Z_{ts},Z_{rp})=\sqrt{V_{\varepsilon,tt}V_{\varepsilon,ss}V_{\varepsilon,rr}V_{\varepsilon,pp}}Cov(\mathscr{Z}_{ts},\mathscr{Z}_{rp})$. Moreover, $Z_{n}^{*}\Rightarrow Z^{*}=G^{\prime}V_{\varepsilon}^{-1}ZV_{\varepsilon}^{-1}G$ and $\bar{Z}_{n}:=Z_{n}-\mathcal{T}_{F,V_{\varepsilon}}(Z_{n})\Rightarrow\bar{Z}$, where $\bar{Z}=Z-\mathcal{T}_{F,V_{\varepsilon}}(Z)=Z-V_{\varepsilon}^{2}diag\left((\boldsymbol{X}^{\prime}\boldsymbol{X})^{-1}\boldsymbol{X}^{\prime}vech(Z^{*})\right)$ (see (B.7)). The distributional limit of $W_{n}$ is given next.
@@
Under Assumptions 1, 2 and 4, 5, 10, as $n\rightarrow\infty$, (a) we have $W_{n}\Rightarrow\bar{W}$, where $vec(\bar{W})\sim N(0,\Omega_{W})$ with $\Omega_{W}=Q_{\beta}\otimes V_{\varepsilon}$, and (b) if additionally $E[w_{i,t}w_{i,r}w_{i,s}]=0$, for all $t,r,s$ and $i$, then $Z$ and $\bar{W}$ are independent.
We get the following proposition from Lemmas 2 and 8 (see proof at the end of the section).
@@
Under Assumptions 1-2 and 3-8, 10, as $n\rightarrow\infty$ and $T$ is fixed, for $j=1,..,k$:
$\hskip-21.33955pt\sqrt{n}diag(\hat{V}_{\varepsilon}-\tilde{V}_{\varepsilon})\Rightarrow V_{\varepsilon}^{2}(\boldsymbol{X}^{\prime}\boldsymbol{X})^{-1}\boldsymbol{X}^{\prime}vech(Z^{*}),$ (B.20)
$\hskip-21.33955pt\sqrt{n}(\hat{F}_{j}-F_{j})\Rightarrow R_{j}(\bar{W}F^{\prime}+F\bar{W}^{\prime}+\bar{Z})V_{\varepsilon}^{-1}F_{j}+\Lambda_{j}\{[V_{\varepsilon}(\boldsymbol{X}^{\prime}\boldsymbol{X})^{-1}\boldsymbol{X}^{\prime}vech({Z}^{*})]\odot F_{j}\},\qquad\quad$ (B.21)
$\hskip-21.33955pt\sqrt{n}(\hat{F}_{j}\hat{\mathcal{D}}-F_{j})\Rightarrow\frac{1}{\gamma_{j}}(\bar{W}F^{\prime}+F\bar{W}^{\prime}+\bar{Z})V_{\varepsilon}^{-1}F_{j},\qquad\quad$ (B.22)
@@
The joint asymptotic Gaussian distribution of the FA estimators involves the Gaussian matrices $^{*}, Z¯\bar{Z}$ and $\bar{W}$, the former two being symmetric. The asymptotic distribution of $\hat{V}_{\varepsilon}$ involves recentering around $\tilde{V}_{\varepsilon}=\frac{1}{n}\sum_{i=1}^{n}E[\varepsilon_{i}\varepsilon_{i}^{\prime}]$, i.e., the finite-sample average cross-moments of errors, and not $V_{\varepsilon}$. For the asymptotic distribution of any functional that depends on $F$ up to one-to-one transformations of its columns, we can use the Gaussian law of (B.22) involving $\bar{W}$ and $\bar{Z}$ only.
The asymptotic expansions (B.20)-(B.21) characterize explicitly the matrices $C_{1}(\theta)$ and $C_{2}(\theta)$ that appear in Theorem 2 in Anderson and Amemiya (1988). Their derivation is based on an asymptotic normality argument treating $\hat{\theta}$ as a M-estimator. However, neither the asymptotic variance nor a feasible CLT are given in Anderson and Amemiya (1988). We cannot use their results for our empirics.
@@
Finally, by applying the CLT to (B.19), the asymptotic distribution of vector $\hat{\theta}$ is:
$\sqrt{n}(\hat{\theta}-\tilde{\theta})\Rightarrow\frac{1}{2}B_{0}\left(B_{0}^{\prime}J_{0}B_{0}\right)^{-1}B_{0}^{\prime}\left(\frac{\partial vec(\Sigma({\theta}_{0}))}{\partial\theta^{\prime}}\right)^{\prime}\left({V}_{y}^{-1}\otimes{V}_{y}^{-1}\right)vec\left(\bar{W}F^{\prime}+F\bar{W}^{\prime}+Z\right).\quad$ (B.23)
The Gaussian asymptotic distribution in (B.23) matches those in (B.20) and (B.21) written for the components, and its asymptotic variance yields the ‘sandwich formula”. The result in (B.23) is analogue to Theorem 2 in Anderson and Amemiya (1988), for different factor normalization and recentering of the variance estimator.
Proof of Proposition 7: From (B.7), we have the asymptotic expansion: $\sqrt{n}diag(\hat{V}_{\varepsilon}-\tilde{V}_{\varepsilon})=diag(\Psi_{\varepsilon})+o_{p}(1)=V_{\varepsilon}^{2}(\boldsymbol{X}^{\prime}\boldsymbol{X})^{-1}\boldsymbol{X}^{\prime}vech(Z_{n}^{*})+o_{p}(1).$ Moreover, from Proposition 6 (a) and using $\Psi_{y}-\Psi_{\varepsilon}=W_{n}F^{\prime}+FW_{n}^{\prime}+\bar{Z}_{n}$, we have: $\sqrt{n}(\hat{F}_{j}-F_{j})=R_{j}(\Psi_{y}-\Psi_{\varepsilon})V_{\varepsilon}^{-1}F_{j}+\Lambda_{j}\Psi_{\varepsilon}V_{\varepsilon}^{-1}F_{j}+o_{p}(1)=R_{j}(W_{n}F^{\prime}+FW_{n}^{\prime}+\bar{Z}_{n})V_{\varepsilon}^{-1}F_{j}+\Lambda_{j}[diag(\Psi_{\varepsilon})\odot(V_{\varepsilon}^{-1}F_{j})]+o_{p}(1)=R_{j}(W_{n}F^{\prime}+FW_{n}^{\prime}+\bar{Z}_{n})V_{\varepsilon}^{-1}F_{j}+\Lambda_{j}\{[V_{\varepsilon}(\boldsymbol{X}^{\prime}\boldsymbol{X})^{-1}\boldsymbol{X}^{\prime}vech(Z_{n}^{*})]\odot F_{j}\}+o_{p}(1).$ Lemmas 2 and 8 yield (B.20)-(B.21), together with (B.22) from (B.8) since $\Psi_{y}-\Psi_{\varepsilon}\Rightarrow\bar{W}F^{\prime}+F\bar{W}^{\prime}+\bar{Z}$.
B.5.3 Feasible CLT for the FA estimators